# Online Resource Allocation via Static Bundle Pricing

### v1

@@ Appendix B Price Subadditivity
After solving the scaled-down LP of Step 1, let $\mathcal{H}$ denote the set of feasible fractional solutions $x$ attaining value $\textnormal{{FracOpt}}_{\gamma}$. Among these, we select a canonical solution by solving the secondary LP
@@
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{S}}x_{b,v},\qquad w_{e}(x)=\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{S}}x_{b,v}.$
@@
Suppose, for the sake of contradiction, that $\sum_{i=1}^{\ell}v_{i}\leq v$. First, if $v=0$ the inequality trivially holds, thus we can assume that $v>0$. Then, if $\sum_{i=1}^{\ell}v_{i}<v$, since $v_{1},\dots,v_{\ell}$ are crucial or tight values and $v$ is non-tight we can transfer $\epsilon>0$ mass (for sufficiently small $\epsilon$) from each $x_{b_{i},v_{i}}>0$ to $x_{b_{j},v}$ for some choice of buyers $b_{i}\in\mathcal{A}_{T_{i}},\,b_{j}\in\mathcal{A}_{S}$ for $i\in[\ell]$, and remain feasible while increasing the objective; contradiction.
@@
Fix a buyer whose desired set is $S$. Suppose that in the posted menu they purchase bundles $T_{1},\dots,T_{\ell}$ with and
@@
Note that since we always offer $|\mathcal{A}_{S}|$ copies of set $S$ at some price, a copy of $S$ will always be available. Indeed, for the buyer that arrives first this trivially holds and since we will prove that the menu is subadditive they will not purchase bundles of any other desired set, thus the $|\mathcal{A}_{S}|$ are enough for each desired set $S$ (they are dedicated to buyers that desire $S$).
First, if $p(S)=1$ then $p(T_{i})=0$, contradiction. Now notice that $x_{T_{i},p(T_{i})}>0$ and $x_{S,p(S)-1}$ is not tight by the definition of our pricing rules. Thus, we can apply Lemma˜B.1 with $v_{i}=p(T_{i})$ and v=p(S)1v=p(S)-1 and get
@@ Appendix C Analysis of the General Algorithmic Framework
Fix a desired bundle $S$ and write $\mathcal{A}=\mathcal{A}_{S}$ and $w=w_{S}$ for its important value. Recall that $\textnormal{{FracOpt}}_{S}\;=\;\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}\,v$ is the contribution of buyers in $\mathcal{A}$ to the scaled-down fractional optimum $\textnormal{{FracOpt}}_{\gamma}$, and $x_{S,v}\;=\;\sum_{b\in\mathcal{A}}x_{b,v}$, $q_{S,v}\;=\;\sum_{b\in\mathcal{A}}q_{b,v}$. By Lemma 2.4, the vector $X^{(S)}=(x_{S,v})_{v\in\mathcal{S}}$ consists of a prefix of zeros, possibly a single crucial value $w$, and then a suffix of tight entries $x_{S,v}=q_{S,v}$ for all $v>w$. We now analyze the expected value obtained by the unconstrained algorithm UAlg from buyers in $\mathcal{A}$ under the menu of Step 2.
Case 1: S has no crucial value. Then $x_{S,v}=0$ for $v\leq w$ and $x_{S,v}=q_{S,v}$ for all $v>w$. The menu offers $|\mathcal{A}|$ copies of $S$ at price $w+1$. Since there are at most $|\mathcal{A}|$ buyers wanting $S$, no buyer is ever blocked in UAlg, and buyers with $v\leq w$ never buy at price $w+1$. Thus
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;=\;\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\,v\;=\;\sum_{b\in\mathcal{A}}\sum_{v>w}x_{b,v}\,v\;=\;\textnormal{{FracOpt}}_{S}.$
Case 2: S has a crucial value w and $x_{S,w}\,w\leq\sum_{v>w}x_{S,v}\,v$. The menu is again $|\mathcal{A}|$ copies of $S$ at price $w+1$, so the same argument as in Case 1 gives
@@
Case 3: S has a crucial value w and $x_{S,w}\,w>\sum_{v>w}x_{S,v}\,v$. Now the mass at $w$ dominates. Let $Y$ denote the number of buyers in $\mathcal{A}$ whose value equals $w$; then $Y$ is a sum of independent binary random variables with total mean $q_{S,w}$. We consider two subcases, depending on $x_{S,w}$.
Case 3a: $x_{S,w}\leq 1$. The menu offers: $|\mathcal{A}|$ copies of $S$ at price $w+1$, and one additional copy at price $w$ with probability $p\;\coloneqq\;\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$. We distinguish according to $q_{S,w}$:
  • If $q_{S,w}\leq 1$, then $x_{S,w}/q_{S,w}\geq x_{S,w}$, so $p=x_{S,w}/q_{S,w}$. By Lemma G.3, $\Pr[Y\geq 1]\geq q_{S,w}/2$. Conditioned on including the $w$-priced copy and on $Y\geq 1$, UAlg sells that copy at value $w$, hence the expected contribution of this copy is at least
    $p\cdot\Pr[Y\geq 1]\cdot w\;\geq\;\frac{x_{S,w}}{q_{S,w}}\cdot\frac{q_{S,w}}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
  • If $q_{S,w}>1$, then $x_{S,w}/q_{S,w}<x_{S,w}$, so $p=x_{S,w}$. Lemma G.3 yields $\Pr[Y\geq 1]\geq 1/2$, and thus the expected contribution of the $w$-priced copy is at least
    $p\cdot\Pr[Y\geq 1]\cdot w\;\geq\;x_{S,w}\cdot\frac{1}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
@@
Case 3b: $x_{S,w}>1$. The menu offers: $\lfloor x_{S,w}\rfloor$ copies of $S$ at price $w$, and $|\mathcal{A}|$ copies at price $w+1$. Lemma G.3 with $\mu=q_{S,w}$ gives
$\Pr\bigl[Y\geq q_{S,w}/2\bigr]\;\geq\;\frac{1}{2},$
@@
$\Pr\bigl[Y\geq x_{S,w}/2\bigr]\;\geq\;\Pr\bigl[Y\geq q_{S,w}/2\bigr]\;\geq\;\frac{1}{2}.$
@@
Fix a desired set $S$ and let $\mathcal{A}=\mathcal{A}_{S}$. We analyze the contribution to the load on a fixed item $e\in S$ coming from buyers in $\mathcal{A}$. As before, we assume that buyers in $\mathcal{A}$ only purchase menu entries corresponding to the bundle $S$; the general case follows from price subadditivity (Appendix B).
For $b\in\mathcal{A}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on item $e$, and set $L_{S}(e)=\sum_{b\in\mathcal{A}}L_{b}(e)$. We inspect each of the menu cases of Step 2.
Case 1: $|\mathcal{A}|$ copies at price $w+1$.
Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{S}(e)=\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]=\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 2: one copy at price $w$ w.p. $x_{S,w}/q_{S,w}$, and $|\mathcal{A}|$ copies at $w+1$.
A buyer contributes to $L_{S}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
$L_{S}(e)\;\leq\;\mathbbm{1}{\left\{\text{$w$-copy is offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b\,:\,v_{b}=w\right\}}\;+\;\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\frac{x_{S,w}}{q_{S,w}}\cdot q_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=x_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 3: one copy at price $w$ w.p. $x_{S,w}$, and $|\mathcal{A}|$ copies at $w+1$.
Exactly as above,
$L_{S}(e)\leq\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq x_{S,w}+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
Case 4: $\lfloor x_{S,w}\rfloor$ copies at price $w$ and $|\mathcal{A}|$ copies at $w+1$.
In this case,
$L_{S}(e)\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{A}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{A}}\sum_{v\in\mathcal{S}}x_{b,v}.$
@@
For Case 2 we will use standard properties from of negative dependence (see Section 2 in [DubhashiRanjan1996]), first set $A_{b}=\mathbbm{1}{\left\{v_{b}=w\right\}}$, $B_{b}=\mathbbm{1}{\left\{v_{b}>w\right\}}$, and $O=\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}$. Then $A_{b}+B_{b}\leq 1$, so $\{A_{b},B_{b}\}$ is NA. By independent union, $\{A_{b},B_{b}\}_{b\in\mathcal{A}}$ is NA. Applying the non-decreasing map $(A_{b})_{b}\mapsto\mathbbm{1}{\left\{\sum_{b}A_{b}\geq 1\right\}}$ gives $\{\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\{B_{b}\}_{b}\}$ is NA. Since $O$ is independent, adding $O$ preserves NA. Applying the non-decreasing map $(e,f)\mapsto e\cdot f$ yields
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{A}}\Big\}\ \text{is NA, done.}$
@@ Appendix D Graph Setting
The proof is similar to the single-minded case. First, fix a desired $(s,t)$ pair and let $\mathcal{A}=\{b:(s_{b},t_{b})=(s,t)\}$. We analyze the contribution to the load on a fixed edge $e$ coming from buyers in $\mathcal{A}$. As before, we assume that buyers in $\mathcal{A}$ only purchase menu entries corresponding to the bundle $(s,t)$; the general case follows from price subadditivity (Appendix D.2).
For $b\in\mathcal{A}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on edge $e$, and set $L_{(s,t)}(e)=\sum_{b\in\mathcal{A}}L_{b}(e)$. We also define $\mathcal{P}_{s,t}^{e}=\{\,p\in\mathcal{P}_{s,t}:e\in p\,\}$, the set of $(s,t)$ paths that contain edge $e$. Lastly, let $x_{>v,p}=\sum_{b\in\mathcal{A}}\sum_{v^{\prime}>v}x_{b,v^{\prime},p}$ and define $x_{>v},\,q_{>v}$ similarly. We inspect each of the menu cases of Step 2.
Case 1: $|\mathcal{A}|$ copies at price $w+1$.
Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{(s,t)}(e)=\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]=q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}\leq\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 2: one copy at price $w$ w.p. $x_{w}/q_{w}$, and $|\mathcal{A}|$ copies at $w+1$.
A buyer contributes to $L_{(s,t)}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
@@
$\;+\;\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq\frac{x_{w}}{q_{w}}\cdot q_{w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}$
$=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{w,p}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}=\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 3: one copy at price $w$ w.p. $x_{w}$, and $|\mathcal{A}|$ copies at $w+1$.
Exactly as above,
@@
$+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq x_{w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{w,p}}{x_{w}}+q_{>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{>w,p}}{x_{>w}}$
$=\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{w,p}+\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{>w,p}=\sum_{b\in\mathcal{A}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{S}}x_{b,v,p}$
Case 4: $\lfloor x_{w}\rfloor$ copies at price $w$ and $|\mathcal{A}|$ copies at $w+1$.
In this case,
$L_{(s,t)}(e)$ $\leq\sum_{j=1}^{\left\lfloor x_{w}\right\rfloor}\mathbbm{1}{\left\{\text{$j$-th copy routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}$
$+\sum_{b\in\mathcal{A}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$

### v2

@@ Appendix B Price Subadditivity
After solving the scaled-down EA-LP of Step 1, in Section 6, let $\mathcal{H}$ denote the set of feasible fractional solutions $x$ attaining value $\textnormal{{FracOpt}}_{\gamma}$. Among these, we select a canonical solution by solving the secondary LP
@@
$\textsc{FracWeight}(x)=\sum_{b}\sum_{v\in\mathcal{U}}x_{b,v},\qquad w_{e}(x)=\sum_{b:e\in S_{b}}\sum_{v\in\mathcal{U}}x_{b,v}.$
@@
Suppose, for the sake of contradiction, that $\sum_{i=1}^{\ell}v_{i}\leq v$. First, if $v=0$ the inequality trivially holds, thus we can assume that $v>0$. Then, if $\sum_{i=1}^{\ell}v_{i}<v$, since $v_{1},\dots,v_{\ell}$ are crucial or tight values and $v$ is non-tight we can decrease $\epsilon>0$ mass (for sufficiently small $\epsilon$) from each $x_{b_{i},v_{i}}>0$ and increase $x_{b_{j},v}$ by $\epsilon$, for some choice of buyers $b_{i}\in\mathcal{N}_{T_{i}},\,b_{j}\in\mathcal{N}_{S}$ for $i\in[\ell]$. Consequently, we remain feasible while increasing the objective; contradiction.
@@
Fix a buyer whose desired set is $S$. Suppose that in the posted menu they purchase bundles $T_{1},\dots,T_{\ell}$ and
@@
Note that since we always offer $|\mathcal{N}_{S}|$ copies of set $S$ at some price, a copy of $S$ will always be available. Indeed, for the buyer that arrives first this trivially holds and since we will prove that the menu is subadditive they will not purchase bundles of any other desired set, thus the $|\mathcal{N}_{S}|$ are enough for each desired set $S$ (they are dedicated to buyers that desire $S$). We could offer $n$ copies (where $n$ is the total number of buyers) and avoid the above argument but $|\mathcal{N}_{S}|$ copies suffice.
First, if $p(S)=1$ then $p(T_{i})=0$, contradiction. Now notice that $x_{T_{i},p(T_{i})}>0$ and $x_{S,p(S)-1}$ is not tight by the definition of our pricing rules. Thus, we can apply Lemma B.1 with $v_{i}=p(T_{i})$ and v=p(S)1v=p(S)-1 and get
@@ Appendix C Analysis of the General Algorithmic Framework
Fix a desired bundle $S$ and write $w=w_{S}$ for its important value. Let $\textnormal{{FracOpt}}_{S}\;=\;\sum_{b\in\mathcal{N}_{S}}\sum_{v\in\mathcal{U}}x_{b,v}\,v$ be the contribution of buyers in $\mathcal{N}_{S}$ to the scaled-down fractional optimum $\textnormal{{FracOpt}}_{\gamma}$, and $x_{S,v}\;=\;\sum_{b\in\mathcal{N}_{S}}x_{b,v}$, $q_{S,v}\;=\;\sum_{b\in\mathcal{N}_{S}}q_{b,v}$. By Lemma 5.2, the vector $X^{(S)}=(x_{S,v})_{v\in\mathcal{U}}$ consists of a prefix of zeros, possibly a single crucial value $w$, and then a suffix of tight entries $x_{S,v}=q_{S,v}$ for all $v>w$. We now analyze the expected value obtained by the unconstrained algorithm UAlg from buyers in $\mathcal{N}_{S}$ under the menu of Step 2, in Section 6.
Case 1: S has no crucial value. Then $x_{S,v}=0$ for $v\leq w$ and $x_{S,v}=q_{S,v}$ for all $v>w$. The menu offers $|\mathcal{N}_{S}|$ copies of $S$ at price $w+1$. Since there are at most $|\mathcal{N}_{S}|$ buyers wanting $S$, no buyer is ever blocked in UAlg, and buyers with $v\leq w$ never buy at price $w+1$. Thus
$\mathbb{E}\left[\textnormal{{UAlg}}\text{ value from }S\right]\;=\;\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}\,v\;=\;\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}x_{b,v}\,v\;=\;\textnormal{{FracOpt}}_{S}.$
Case 2: S has a crucial value w and $x_{S,w}\,w\leq\sum_{v>w}x_{S,v}\,v$. The menu is again $|\mathcal{N}_{S}|$ copies of $S$ at price $w+1$, so the same argument as in Case 1 gives
@@
Case 3: S has a crucial value w and $x_{S,w}\,w>\sum_{v>w}x_{S,v}\,v$. Now the mass at $w$ dominates. Let $Y$ denote the number of buyers in $\mathcal{N}_{S}$ whose value equals $w$; then $Y$ is a sum of independent binary random variables with total mean $q_{S,w}$. We consider two subcases, depending on $x_{S,w}$.
Case 3a: $x_{S,w}\leq 1$. The menu offers: $|\mathcal{N}_{S}|$ copies of $S$ at price $w+1$, and one additional copy at price $w$ with probability $p\;\coloneqq\;\max\{x_{S,w},\,x_{S,w}/q_{S,w}\}$. We distinguish according to $q_{S,w}$:
  • If $q_{S,w}\leq 1$, then $x_{S,w}/q_{S,w}\geq x_{S,w}$, so $p=x_{S,w}/q_{S,w}$. By Lemma F.4, $\mathbb{P}\!\left[Y\geq 1\right]\geq q_{S,w}/2$. Conditioned on including the $w$-priced copy and on $Y\geq 1$, UAlg sells that copy at value $w$, hence the expected contribution of this copy is at least
    $p\cdot\mathbb{P}\!\left[Y\geq 1\right]\cdot w\;\geq\;\frac{x_{S,w}}{q_{S,w}}\cdot\frac{q_{S,w}}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
  • If $q_{S,w}>1$, then $x_{S,w}/q_{S,w}<x_{S,w}$, so $p=x_{S,w}$. Lemma F.4 yields $\mathbb{P}\!\left[Y\geq 1\right]\geq 1/2$, and thus the expected contribution of the $w$-priced copy is at least
    $p\cdot\mathbb{P}\!\left[Y\geq 1\right]\cdot w\;\geq\;x_{S,w}\cdot\frac{1}{2}\cdot w\;=\;\frac{1}{2}\,x_{S,w}\,w.$
@@
Case 3b: $x_{S,w}>1$. The menu offers: $\lfloor x_{S,w}\rfloor$ copies of $S$ at price $w$, and $|\mathcal{N}_{S}|$ copies at price $w+1$. Lemma F.4 with $\mu=q_{S,w}$ gives
$\mathbb{P}\!\left[Y\geq q_{S,w}/2\right]\;\geq\;\frac{1}{2},$
@@
$\mathbb{P}\!\left[Y\geq x_{S,w}/2\right]\;\geq\;\mathbb{P}\!\left[Y\geq q_{S,w}/2\right]\;\geq\;\frac{1}{2}.$
@@
Fix a desired set $S$. We analyze the contribution to the load on a fixed item $e\in S$ coming from buyers in $\mathcal{N}_{S}$. As before, we assume that buyers in $\mathcal{N}_{S}$ only purchase menu entries corresponding to the bundle $S$; the general case follows from price subadditivity (Appendix B).
For $b\in\mathcal{N}_{S}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on item $e$, and set $L_{S}(e)=\sum_{b\in\mathcal{N}_{S}}L_{b}(e)$. We inspect each of the menu cases of Step 2, in Section 6.
Case 1: $|\mathcal{N}_{S}|$ copies at price $w+1$. Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{S}(e)=\sum_{b\in\mathcal{N}_{S}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]=\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{N}_{S}}\sum_{v\in\mathcal{U}}x_{b,v}.$
Case 2: one copy at price $w$ w.p. $x_{S,w}/q_{S,w}$, and $|\mathcal{N}_{S}|$ copies at $w+1$. A buyer contributes to $L_{S}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
$L_{S}(e)\;\leq\;\mathbbm{1}{\left\{\text{$w$-copy is offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b\,:\,v_{b}=w\right\}}\;+\;\sum_{b\in\mathcal{N}_{S}}\mathbbm{1}{\left\{v_{b}>w\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\frac{x_{S,w}}{q_{S,w}}\cdot q_{S,w}+\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}=x_{S,w}+\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{N}_{S}}\sum_{v\in\mathcal{U}}x_{b,v}.$
Case 3: one copy at price $w$ w.p. $x_{S,w}$, and $|\mathcal{N}_{S}|$ copies at $w+1$. Exactly as above,
$L_{S}(e)\leq\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}+\sum_{b\in\mathcal{N}_{S}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq x_{S,w}+\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}=\sum_{b\in\mathcal{N}_{S}}\sum_{v\in\mathcal{U}}x_{b,v}.$
Case 4: $\lfloor x_{S,w}\rfloor$ copies at price $w$ and $|\mathcal{N}_{S}|$ copies at $w+1$. In this case,
$L_{S}(e)\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{N}_{S}}\mathbbm{1}{\left\{v_{b}>w\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{S}(e)\right]\leq\lfloor x_{S,w}\rfloor+\sum_{b\in\mathcal{N}_{S}}\sum_{v>w}q_{b,v}\leq\sum_{b\in\mathcal{N}_{S}}\sum_{v\in\mathcal{U}}x_{b,v}.$
@@
For Case 2 we will use standard properties of negative dependence (see Section 2 in Dubhashi and Ranjan [1998]), first set $A_{b}=\mathbbm{1}{\left\{v_{b}=w\right\}}$, $B_{b}=\mathbbm{1}{\left\{v_{b}>w\right\}}$, and $O=\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}$. Then $A_{b}+B_{b}\leq 1$, so $\{A_{b},B_{b}\}$ is NA. By independent union, $\{A_{b},B_{b}\}_{b\in\mathcal{N}_{S}}$ is NA. Applying the non-decreasing map $(A_{b})_{b}\mapsto\mathbbm{1}{\left\{\sum_{b}A_{b}\geq 1\right\}}$ gives $\{\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\{B_{b}\}_{b}\}$ is NA. Since $O$ is independent, adding $O$ preserves NA. Applying the non-decreasing map $(e,f)\mapsto e\cdot f$ yields
$\Big\{\mathbbm{1}{\left\{\text{$w$-copy offered}\right\}}\cdot\mathbbm{1}{\left\{\exists b:\,v_{b}=w\right\}},\ \{\mathbbm{1}{\left\{v_{b}>w\right\}}\}_{b\in\mathcal{N}_{S}}\Big\}\ \text{is NA, done.}$
@@ Appendix D Graph Setting
The proof is similar to the single-minded case. First, fix a desired $(s,t)$ pair. We analyze the contribution to the load on a fixed edge $e$ coming from buyers in $\mathcal{N}_{s,t}$. As before, we assume that buyers in $\mathcal{N}_{s,t}$ only purchase menu entries corresponding to the bundle $(s,t)$; the general case follows from price subadditivity (Appendix D.2).
For $b\in\mathcal{N}_{s,t}$ let $L_{b}(e)$ be the indicator that buyer $b$ contributes to the load on edge $e$, and set $L_{(s,t)}(e)=\sum_{b\in\mathcal{N}_{s,t}}L_{b}(e)$. Recall that $\mathcal{P}_{s,t}^{e}$ denotes the set of $(s,t)$ paths that contain edge $e$. Lastly, let $x_{(s,t),\geq v,p}=\sum_{b\in\mathcal{N}_{s,t}}\sum_{v^{\prime}\geq v}x_{b,v^{\prime},p}$ and $x_{(s,t),\geq v}=\sum_{b\in\mathcal{N}_{s,t}}\sum_{v^{\prime}\geq v}\sum_{p\in\mathcal{P}_{s,t}}x_{b,v^{\prime},p}$. We inspect each of the menu cases of Step 2, in Section 6.
Case 1: $|\mathcal{N}_{s,t}|$ copies at price $w+1$. Here a buyer accepts if and only if its realized value is $v>w$. Thus
$L_{(s,t)}(e)=\sum_{b\in\mathcal{N}_{s,t}}\mathbbm{1}{\left\{v_{b}>w\right\}}\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]\leq q_{(s,t),>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}\leq\sum_{p\in\mathcal{P}_{s,t}^{e}}x_{(s,t),\geq w,p}\leq\sum_{b\in\mathcal{N}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{U}}x_{b,v,p}$
Case 2: one copy at price $w$ w.p. $x_{(s,t),w}/q_{(s,t),w}$, and $|\mathcal{N}_{s,t}|$ copies at $w+1$. A buyer contributes to $L_{(s,t)}(e)$ either by (i) obtaining the (possibly offered) $w$-copy, or (ii) having value $v>w$. Since at most one $w$-copy can be claimed,
@@
$\;+\;\sum_{b\in\mathcal{N}_{s,t}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{$b$ routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}.$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq\frac{x_{(s,t),w}}{q_{(s,t),w}}\cdot q_{(s,t),w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}+q_{(s,t),>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}$
$=\sum_{b\in\mathcal{N}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{U}}x_{b,v,p}$
Case 3: one copy at price $w$ w.p. $x_{(s,t),w}$, and $|\mathcal{N}_{s,t}|$ copies at $w+1$. Exactly as above,
@@
$+\sum_{b\in\mathcal{N}_{s,t}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$
@@
$\mathbb{E}_{M,I}\left[L_{(s,t)}(e)\right]$ $\leq x_{(s,t),w}\cdot\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}+q_{(s,t),>w}\sum_{p\in\mathcal{P}_{s,t}^{e}}\frac{x_{(s,t),\geq w,p}}{x_{(s,t),\geq w}}$
$=\sum_{b\in\mathcal{N}_{s,t}}\sum_{p\in\mathcal{P}_{s,t}^{e}}\sum_{v\in\mathcal{U}}x_{b,v,p}$
Case 4: $\lfloor x_{(s,t),w}\rfloor$ copies at price $w$ and $|\mathcal{N}_{s,t}|$ copies at $w+1$. In this case,
$L_{(s,t)}(e)$ $\leq\sum_{j=1}^{\left\lfloor x_{(s,t),w}\right\rfloor}\mathbbm{1}{\left\{\text{$j$-th copy routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}}$
$+\sum_{b\in\mathcal{N}_{s,t}}\mathbbm{1}{\left\{v_{b}>w\right\}}\cdot\mathbbm{1}{\left\{\text{buyer routed through a path in $\mathcal{P}_{s,t}^{e}$}\right\}},$